Small HW on Tariffs g Diane Monaco Market for Sugar Based Ethanol (SBE) Supply* Price Of SBE D E P IMPORTS B Port C
Answer #1

Pworld $3.40 Pleweshe = $3.90 Fisheriff $.15 5 Presid, tariff = $3.55 At 3 of domestic da domestic 34 At domestic 20 At Portawith trade A AE $3.40 A domestica s Od domestic 34 impost = 34-5 29 unit C.5 = 1 (8-3.40) x 34 $78.2 10 P.S - / (3.40 - 2.10)Camlin Page Date Dlintercept) = $19.20 s(intercept) & 6.80 2. w/o trade At Pa = $11.80 134.75 Casya ut (19.20 - 11.50),35 P.SDWL = (10.90-10.20) (14-87} { (10.90-10.2.0) (25 – 27) Due overproduction Due to under consumption 5 12-1 + 2.8 .8